The canola oil industry is perfectly competitive. Every producer has the following total cost function: LTC...

  1. The canola oil industry is perfectly competitive. Every producer has the following total cost function: LTC = 2Q3 – 15Q2 + 40Q, where Q is measured in tons of canola oil. The corresponding marginal cost function is given by LMC = 6Q2 – 30Q + 40.

a. In long-run equilibrium, how much will each firm produce?

b. What is the long-run equilibrium price?

c. Suppose that the market demand for canola oil is given by Q = 999 – 0.25P. At the long-run equilibrium price, how many tons of canola oil will consumers demand?
Answer #1

a LTC= 203 - 1502 +400 AC = 202 150+40 DAC 40-15= 0 (0=15/9 = 3.5 ACF 2 (3.25² - 15(3.8) +40 – 98. 185-58.25 +yo [AC = 11.88/